Given the function f(x) = - 4x find the AVERAGE RATE of CHANGE of f from -4 to 1
Answer by stanbon(75887)

Given the function f(x) = - 4x find the AVERAGE RATE of CHANGE of f from -4 to 1
-----------
Average = [f(1)-f(-4)]/[1--4]
Average = [-4 - 16] / [1+4] = -20/5 = -4
Notice that the answer is the slope of y = -4x
=======================
Cheers,
Stan H.